OpenStudy (anonymous):

Find an exact value. cos 15° Options: A) sqrt 6 + sqrt 2 / 4 B) - sqrt2 + sqrt 6 / 4 C) - sqrt2 - sqrt 6 / 4 D) - sqrt6 + 1 / 4

OpenStudy (anonymous):

A) \[ \frac{ \sqrt{6}+\sqrt{2} }{ 4 } \] B) \[\frac{ -\sqrt{2} +\sqrt{6} }{ 4 }\] C) \[\frac{ -\sqrt{2} - \sqrt{6} }{ 4 }\] D) \[\frac{ -\sqrt{6} +1}{ 4 }\]

OpenStudy (amoodarya):

cos(a-b)=cosa cosb + sina sin b you know? if ok cos (45-30)= cos45 cos 30 +sin 45 sin 30

OpenStudy (anonymous):

where did you get all the numebrs?

OpenStudy (amoodarya):

 I dont get what you mean !

OpenStudy (anonymous):

But its for cos 15° ?

OpenStudy (amoodarya):

It is an unknown angle you have to make it bu known angle

OpenStudy (amoodarya):

you know that formula? if yeah put the numbers in it you will have answer (a)

OpenStudy (anonymous):

I think I got it. Can you help with another?

OpenStudy (klimenkov):

Also, you can use \(\cos^2\frac x 2=\frac{1+\cos x}2\).

OpenStudy (anonymous):

@klimenkov Can you help with this one, pelase? Write the expression as either the sine, cosine, or tangent of a single angle. \[\cos(\frac{ \pi }{ 5 }) \cos(\frac{ \pi }{ 7 }) + \sin (\frac{ \pi }{ 5}) \sin (\frac{ \pi }{ 7 })\]

OpenStudy (amoodarya):

as "klimenkov " says you can use that formula shuch the way i draw